Using the information in the table below, calculate the postoperative death rate at Community Hospital for...

  1. Using the information in the table below, calculate the postoperative death rate at Community Hospital for the Surgery Service for the month of July.

Express as a percent and round to two decimal places.

  • Discharges                                                            309
  • Deaths:                                                      005
  • Within 10 days of surgery                   003
  • After 10 days of surgery                     002
  • Number of operations                              290
  • Number of patients operated on              284

please show your steps
Answer #1

Postoperative death rate.

Postoperative death rate= Number of deaths within 10 days after surgery×100/Total number of patients operated on for the period.

Number of deaths within 10 days after surgery=3

number of patients operated on for the period = 284

Postoperative death rate for the month of July= 3×100/ 284

=300/284 = 1.056.

Postoperative death rate for the month of July: 1.05%